OpenStudy (anonymous):

radical negative 36 times radical negative eighteen?

TheSmartOne (thesmartone):

\[\sqrt{a}\times \sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

negative

TheSmartOne (thesmartone):

\[\sqrt{-36}\times \sqrt{-18}=\sqrt{-18\times-36}=\sqrt{648}=18\sqrt{2}\]